OpenStudy (anonymous):

Use the substitution method to solve the system of equations. Enter your answer as an ordered pair. 14x - 2y = 78 2x - 2y = 6

OpenStudy (anonymous):

put x=y+3 14y+42-2y=78 y=3 x=6 (6,3)